What is the equivalent capacitance between points
Answer #1

20) C

first two are in series = 0.5 uF

now consider the parallel = 0.5 + 1 = 1.5

Now, we have 1.5 and 1 in series

c= 0.6 uF

21) C

c di = do f / (do - f) = 8*10/(8-10) = -40 cm; m = -di/do = -40/8 = -5

22) B

n1 sini = n2 sinr (Simply use Snell's law)

23) d

Kirchhof's law => Sum of currents entering a point = sum of currents leaving